Find the Net Present Value of a project’s cash flow which costs $75,000 and provides cash flows of $9,200 per year for six years with annual escalations of 2.5% per year. Assume an 8% annual rate of return.
Answer #1

Year CF Year ICF 0 $(75,000.00) O |-75000 1 $ 9,200.00 1 19200 2 $ 9,430.00 2 I=B3*1.025 3 $ 9,665.75 3 =B4*1.025 4 $ 9,907.3